### v1

# A note on maximal plane subgraphs of the complete twisted graph containing perfect matchings

Abstract
The twisted graph $T_{n}$ is a drawing of the complete graph with $n$ vertices $v_{1},v_{2},\ldots,v_{n}$ in which two edges $v_{i}v_{j}$ ($i<j$) and $v_{s}v_{t}$ ($s<t$) cross if and only if $i<s<t<j$ or $s<i<j<t$. We show that for any maximal plane subgraphs $S$ and $R$ of $T_{n}$, each containing at least one perfect matching, there is a sequence $S=F_{0},F_{1},\ldots,F_{m}=R$ of maximal plane subgrahs of $T_{n}$, also containing perfect matchings, such that for $i=0,1,\ldots,m-1$, $F_{i+1}$ can be obtained from $F_{i}$ by a single edge exchange.
Keywords. Twisted Graph, Max Graph, Matching Graph.
1 Introduction
Consider a set of points $P$ in the euclidean plane. A topological graph with vertex set $P$ is a simple graph with edges drawn in the plane as Jordan curves in such a way that any two edges have at most one point in common. A geometric graph is a topological graph in which all edges are straight line segments. A geometric graph is a convex geometric graph if the set of vertices is in convex position.
Two topological graphs $G$ and $G^{\prime}$ with vertex sets $P$ and $P^{\prime}$, respectively, are weakly isomorphic if there is a bijection $\alpha:P\rightarrow P^{\prime}$ such that $\alpha\left(u\right)\alpha\left(v\right)\in E\left(G^{\prime}\right)$ if and only if $uv\in E\left(G\right)$ and two edges $\alpha\left(x\right)\alpha\left(y\right)$ and $\alpha\left(u\right)\alpha\left(v\right)$ of $G^{\prime}$ intersect if and only if $xy$ and $uv$ intersect in $G$.
The twisted graph $T_{n}$ is a complete topological graph with $n$ vertices $v_{1},v_{2},\allowbreak\ldots,v_{n}$ in which two edges $v_{i}v_{j}$ ($i<j$) and $v_{s}v_{t}$ ($s<t$) cross each other if and only if $i<s<t<j$ or $s<i<j<t$. See Fig. 1
Figure 1: The twisted graph $T_{6}$.
The graphs $T_{n}$, $n\geq 5$ were introduced by Harborth and Mengersen [1] as examples of complete topological graphs containing no subgraphs weakly isomorphic to the complete convex geometric graph $C_{5}$ on 5 vertices. Later, Pach et al. [5] proved that every complete topological graph with $n$ vertices contains a complete subgraph with $m\geq c\log^{1/8}n$ vertices which is weakly isomorphic to either the complete convex geometric graph $C_{m}$ or to the twisted graph $T_{m}$.
Let $F_{n}$ be a complete topological graph with $n$ vertices. A plane subgraph $G$ of $F_{n}$ is a maximal plane subgraph of $F_{n}$ if for each edge $uv$ of $F_{n}$, not in $G$, there is an edge $xy$ of $G$ that intersects $uv$. The max graph $MP\left(F_{n}\right)$ of $F_{n}$ is the abstract graph whose vertices are the maximal plane subgraphs of $F_{n}$ in which two graphs $G$ and $H$ are adjacent if there are edges $g$ of $G$ and $h$ of $H$ such that $H=\left(G-g\right)+h$. In the case where $F_{n}$ is a complete geometric graph with vertex set $P$, the graph $MP\left(F_{n}\right)$ is called the graph of triangulations of $P$ and is denoted by $\mathcal{T}(P)$.
Lawson [3] proved that for each set of points $P$ in general position in the plane, the graph $\mathcal{T}(P)$ is connected and Houle et al [2] showed that if $P$ has an even number of points, the subgraph $\mathcal{T_{M}}(P)$ of $\mathcal{T}(P)$, induced by those triangulations of $P$ that contain perfect matchings, is also connected.
Later, Omaña-Pulido and Rivera-Campo [4] showed that $MP\left(T_{n}\right)$ is always connected. In this note we use some of the ideas of Houle et al [2] and of Omaña-Pulido and Rivera-Campo [4] to show that for any maximal plane subgraphs $R$ and $S$ of $T_{2m}$, each one containing at least one perfect matching, there is a sequence $R=F_{0},F_{1},\ldots,F_{t}=S$ of maximal plane subgrahs of $T_{2m}$, also containing perfect matchings, such that for $i=0,1,\ldots,t-1$, $F_{i+1}$ can be obtained from $F_{i}$ by a single edge exchange.
Throughout the paper we denote by $v_{1},v_{2},\ldots,v_{n}$ the vertices of $T_{n}$ from left to right as in Fig. 1.
2 Preliminary results
The matching graph $\mathcal{M}(T_{2m})$ of $T_{2m}$ is the abstract graph in which the vertices are the plane perfect matchings of $T_{2m}$ and two matchings $L$ and $N$ are adjacent if the symmetric difference $L\bigtriangleup N$ is a plane cycle with four edges. Omaña-Pulido and Rivera-Campo [4] proved that $\mathcal{M}(T_{2m})$ is always connected.
Theorem 1.
[4] For each positive integer $m$, the graph $\mathcal{M}\left(T_{2m}\right)$ is connected.
Let $n\geq 2$ be an integer and $R$ be a maximal plane subgraph of $T_{n}$.
For $i=1,2,\ldots,n-1$ let $N_{R}^{+}(v_{i})=\{v_{j}:v_{i}v_{j}\in E(R)\text{ with }j>i\}$, $d_{R}^{+}(v_{i})=|N_{R}^{+}(v_{i})|$ and $max(v_{i})=max\{j:v_{j}\in N_{R}^{+}(v_{i})\}$. For any pair $R$, $Q$ of maximal plane subgraphs of $T_{n}$ let
$k(R,Q)=\begin{cases}max\{t:d_{R}^{+}(v_{t})=d_{Q}^{+}(v_{t})\}&\quad\text{ if }d_{R}^{+}(v_{1})=d_{Q}^{+}(v_{1})\\ 0&\quad\text{ otherwise }\\ \end{cases}$
and let $m(R,Q)=n-1-k(R,Q)$.
Let $F$ be a set of pairwise non-crossing edges of $T_{n}$. We denote by $MP(T_{n},F)$ the subgraph of $MP(T_{n})$ induced by the maximal planar subgraphs of $T_{n}$ that contain all edges in $F$. The following result will be used, together with Theorem 1, in the proof of our main result.
Theorem 2.
For each integer $n\geq 2$ and each set $F$ of pairwise non-crossing edges of $T_{n}$, the graph $MP(T_{n},F)$ is connected.
Proof.
Let $R$ an $Q$ be maximal planar subgraphs of $T_{n}$ with $F\subset E(R)\cap E(Q)$. We shall proof by Induction on $m(R,Q)$ that $MP(T_{n},F)$ contains a path joining $R$ and $Q$.
If $m(R,Q)=0$, then $k(R,Q)=n-1$ which implies $R=Q$. We proceed assuming that for some integer $t\geq 0$ the result holds whenever $R^{\prime}$ and $Q^{\prime}$ are maximal planar subgraphs of $T_{n}$ with $F\subset E(R^{\prime})\cap E(Q^{\prime})$ and $m(R^{\prime},Q^{\prime})\leq t$.
Let Let $R$ an $Q$ be maximal planar subgraphs of $T_{n}$ with $F\subset E(R)\cap E(Q)$ and $m(R^{\prime},Q^{\prime})=t+1$. Without loss of generality we assume $F=E(R^{\prime})\cap E(Q^{\prime})$.
Case $k(R,Q)=0$.
Also without loss of generality we assume $d_{R}^{+}(v_{1})>d_{Q}^{+}(v_{1})$. Let
[…]
Figure 2: Graphs $R$ (above) and $S_{1}$ (bellow).
We claim $S_{1}$ is a maximal plane subgraph of $T_{n}$ with $F\subset S_{1}$ and therefore $S_{1}$ is adjacent to $R$ in $MP(T_{n},F)$.
If $d_{S_{1}}^{+}(v_{1})=d_{Q}^{+}(v_{1})$, then $k(S_{1},Q)\geq 1$ and $m(S_{1},Q)\leq t$. By the Induction Hypothesis $MP(T_{n},F)$ contains a path joining $S_{1}$ and $R$ which, together with the edge $RS_{1}$ gives a path in $MP(T_{n},F)$ that joins $R$ and $Q$.
If $d_{S_{1}}^{+}(v_{1})>d_{Q}^{+}(v_{1})$ we iterate the process defining, one at a time, maximal plane subgraphs $S_{2},S_{3},\ldots,S_{l}$ of $T_{n}$, where $l=max_{R}(v_{1})-max_{Q}(v_{1})$ and for $i=1,2,\ldots,l-1$
$S_{i+1}=(S_{i}-v_{1}v_{max_{S_{i}}(v_{1})})+v_{2}v_{max_{S_{i}}(v_{1})-1}$
Since $F=E(R)\cap E(Q)$ and $v_{1}v_{i}\in E(R)\setminus E(Q)$ for $i=max_{Q}(v_{1}),max_{Q}(v_{1})+1,\ldots,l$, $F\subset E(S_{i})$ for $i=1,2,\ldots,l$. Therefore $R,S_{1},S_{2},\ldots,S_{l}$ is a path in $MP(T_{n},F)$ joining $R$ and $S_{l}$.
Notice that $d_{S{i+1}}(v_{1})=d_{S{i}}(v_{1})-1$ for $i=1,2,\ldots,l-1$, therefore $d_{l}(v_{1})=d_{R}(v_{1})$. Then $k(S_{l},Q)\geq 1$ and $m(S_{l},Q)\leq t$. Again by the Induction Hypothesis, $MP(T_{n},F)$ contains a path joining $S_{l}$ and $Q$ which, together with the path $R,S_{1},S_{2},\ldots,S_{l}$ forms a path in $MP(T_{n},F)$ that joins $R$ and $Q$.
Case $k=k(R,Q)\geq 1$.
In this case $d_{R}(v_{i})=d_{Q}(v_{1})$ for $i=1,2,\ldots$ and $d_{R}(v_{k+1})\neq d_{Q}(v_{k+1})$. As before, without loss of generality we assume $d_{R}(v_{k+1})>d_{Q}(v_{k+1})$. We claim that if we change $v_{1}$ by $v_{k+1}$ in the argument in Case $k(R,Q)=0$, we obtain a path $R,S^{\prime}_{1},S^{\prime}_{2},\ldots,S^{\prime}_{l^{\prime}}$ in $MP(T_{n},F)$ that joins $R$ and $S^{\prime}_{l^{\prime}}$, where $l^{\prime}=max_{R}(v_{k+1})-max_{Q}(v_{k+1})$ and $d_{S^{\prime}_{l^{\prime}}}(v_{k+1})=d_{Q}(v_{k+1})$.
Then $k(S^{\prime}_{l^{\prime}},Q)\geq k(R,Q)+1$ and $m(S^{\prime}_{l^{\prime}},Q)\leq t$. By the Induction Hypothesis $MP(T_{n},F)$ contains a path joining $S^{\prime}_{l^{\prime}}$ and $Q$ which, together with the path $R,S^{\prime}_{1},S^{\prime}_{2},\ldots,S^{\prime}_{l^{\prime}}$ forms a path in $MP(T_{n},F)$ that join $R$ and $Q$.
3 Main result
As shown in Fig. 3, not all maximal plane subgraphs of $T_{2m}$ contain perfect matchings. Let $MP_{\cal{M}}(T_{2m})$ denote the subgraph of $MP(T_{2m})$ induced by the set of maximal plane subgraphs of $T_{2m}$ that contain at least one perfect matching.
Figure 3: A maximal plane subgraph ot $T_{6}$ with no perfect matchings.
Theorem 3.
For each positive integer $m$, the graph $MP_{\cal{M}}(T_{2m})$ is connected.
Proof.
Let $S$ and $R$ be maximal plane subgraphs of $T_{2m}$, each containing at least one perfect matching, and let $M(S)$ and $M(R)$ be perfect matchings of $S$ and $R$, respectively. By Theorem 1, the graph $\mathcal{M}\left(T_{2m}\right)$ is connected and therefore contains a path $M(S)=M_{0},M_{1},\ldots,M_{k}=M(R)$ joining $M(S)$ and $M(R)$.
By definition, for $i=1,2,\ldots,k$, $M_{i}$ is obtained from $M_{i-1}$ by exchanging the edges of an alternating non-crossing cycle with four edges. Therefore, for $i=1,2,\ldots,k$, $M_{i-1}\cup M_{i}$ is a set of pairwise non-crossing edges of $T_{2m}$ that can be extended to a maximal plane subgraph $S_{i}$ of $T_{2m}$.
Let $S_{0}=S$ and $S_{k+1}=R$. Notice that for $i=0,1,\ldots,k$, $S_{i}$ and $S_{i+1}$ are maximal plane graphs of $T_{2m}$ that contain $M_{i}$. By Theorem 2, the graph $MP(T_{n},M_{i})$ contains a path joining $S_{i}$ and $S_{i+1}$. Clearly $MP(T_{n},M_{i})$ is a subgraph of $MP_{\cal{M}}(T_{2m})$ for each $i=0,1,\ldots,k$. Therefore the graph $MP_{\cal{M}}(T_{2m})$ contains a path joining $S$ and $R$. ∎