### v2

# A note on maximal plane subgraphs of the complete twisted graph containing perfect matchings

Abstract
The twisted graph $T_{n}$ is a drawing of the complete graph with $n$ vertices $v_{1},v_{2},\ldots,v_{n}$ in which two edges $v_{i}v_{j}$ ($i<j$) and $v_{s}v_{t}$ ($s<t$) cross if and only if $i<s<t<j$ or $s<i<j<t$. We show that for any maximal plane subgraphs $S$ and $R$ of $T_{n}$, each containing at least one perfect matching, there is a sequence $S=F_{0},F_{1},\ldots,F_{t}=R$ of maximal plane subgraphs of $T_{n}$, also containing perfect matchings, such that for $i=0,1,\ldots,t-1$, $F_{i+1}$ can be obtained from $F_{i}$ by a single edge exchange i.e. $F_{i+1}=((F_{i}-e)+f)$, where $e$ is an edge of $F_{i}$ not in $F_{i+1}$ and $f$ is an edge of $F_{i+1}$ not in $F_{i}$.
Keywords. Twisted Graph, Max Graph, Matching Graph.
1 Introduction
Consider a set of points $P$ in the Euclidean plane. A topological graph with vertex set $P$ is a simple graph whose edges are drawn in the plane as Jordan curves in such a way that any two edges have at most one point in common and any pair of non-incident edges with a point in common must cross at that point. A topological graph $G$ is a geometric graph if the vertices of $G$ are in general position and all edges of $G$ are straight line segments. A geometric graph is a convex geometric graph if the set of vertices is in convex position.
Two topological graphs $G$ and $G^{\prime}$ with vertex sets $P$ and $P^{\prime}$ and edge sets $E$ and $E^{\prime}$, respectively, are said to be weakly isomorphic if there is a graph isomorphism $\alpha:P\rightarrow P^{\prime}$ such that for all $x,y,u,v\in P$, two edges $xy$ and $uv$ cross each other in $G$ if and only if the corresponding edges $\alpha\left(x\right)\alpha\left(y\right)$ and $\alpha\left(u\right)\alpha\left(v\right)$ cross each other in $G^{\prime}$.
The twisted graph $T_{n}$ is a complete topological graph with $n$ vertices $v_{1},v_{2},\allowbreak\ldots,v_{n}$ in which two edges $v_{i}v_{j}$ ($i<j$) and $v_{s}v_{t}$ ($s<t$) cross each other if and only either $i<s<t<j$ or $s<i<j<t$. See Fig. 1
Figure 1: The twisted graph $T_{6}$.
Harborth and Mengersen [2] introduced twisted graphs as examples of complete topological graphs containing no subgraphs weakly isomorphic to the complete convex geometric graph $C_{5}$ on 5 vertices. Pach et al. [6] proved that if $G$ is a complete topological graph with $n$ vertices, then $G$ contains a complete subgraph with $m\geq c\log^{1/8}n$ vertices which is weakly isomorphic to either the complete convex geometric graph $C_{m}$ or to the twisted graph $T_{m}$. Recently, Suk and Zeng [7] improved this bound to $\log^{1/4-o(1)}n$. Twisted graphs have been generalized by Aichholzer et al. [1].
Given a complete topological graph $F_{n}$ with $n$ vertices. A plane subgraph $G$ of $F_{n}$ is a maximal plane subgraph of $F_{n}$ if for each edge $uv$ of $F_{n}$, not in $G$, there is an edge $xy$ of $G$ that crosses $uv$. The max graph $MP\left(F_{n}\right)$ of $F_{n}$ is the abstract graph whose vertices are the maximal plane subgraphs of $F_{n}$ in which two graphs $G$ and $H$ are adjacent if there are edges $g$ of $G$ and $h$ of $H$ such that $H=\left(G-g\right)+h$. In the case where $F_{n}$ is a complete geometric graph with vertex set $R$, the graph $MP\left(F_{n}\right)$ is called the graph of triangulations of $R$ and is denoted by $\mathcal{T}(R)$. Lawson [4] proved that for each set of points $R$ in general position in the plane, the graph $\mathcal{T}(R)$ is connected. Later, Omaña-Pulido and Rivera-Campo [5] showed that $MP\left(T_{n}\right)$ is also connected.
Houle et al. [3] showed that if $R$ has an even number of points, the subgraph $\mathcal{T_{M}}(R)$ of $\mathcal{T}(R)$, induced by those triangulations of $R$ that contain perfect matchings, is also connected. In this note we use some of the ideas of Houle et al. [3] and of Omaña-Pulido and Rivera-Campo [5] to show that for any maximal plane subgraphs $R$ and $S$ of $T_{2m}$, each one containing at least one perfect matching, there is a sequence $R=F_{0},F_{1},\ldots,F_{t}=S$ of maximal plane subgraphs of $T_{2m}$, also containing perfect matchings, such that for $i=0,1,\ldots,t-1$, $F_{i+1}$ can be obtained from $F_{i}$ by a single edge exchange.
Throughout the paper we assume $T_{n}$ is represented as in Fig. 1 with its vertices denoted by $v_{1},v_{2},\ldots,v_{n}$ from left to right.
2 Preliminary results
The matching graph $\mathcal{M}(T_{2m})$ of $T_{2m}$ is the abstract graph in which the vertices are the plane perfect matchings of $T_{2m}$ and two matchings $L$ and $N$ are adjacent if the symmetric difference $L\bigtriangleup N$ is a plane cycle with four edges. Omaña-Pulido and Rivera-Campo [5] proved that $\mathcal{M}(T_{2m})$ is always connected.
Theorem 1.
[5] For each positive integer $m$, the graph $\mathcal{M}\left(T_{2m}\right)$ is connected.
Let $n\geq 3$ be an integer and $R$ be a maximal plane subgraph of $T_{n}$. For $i=1,2,\ldots,n-1$, let $N_{R}^{+}(v_{i})=\{v_{j}:v_{i}v_{j}\in E(R)\text{ with }j>i\}$, $d_{R}^{+}(v_{i})=|N_{R}^{+}(v_{i})|$, $min_{R}(v_{i})=min\{j:v_{j}\in N_{R}^{+}(v_{i})\}$ and $max_{R}(v_{i})=max\{j:v_{j}\in N_{R}^{+}(v_{i})\}$.
Lemma 2.
Let $n\geq 3$ be an integer and $R$ be a maximal plane subgraph of $T_{n}$. Then:
i) All edges $v_{1}v_{2},v_{1}v_{3},v_{n-2}v_{n}$ and $v_{n-1}v_{n}$ are edges of $R$,
ii) For $i=1,2,\ldots,n-2$, $min_{R}(v_{i+1})=max_{R}(v_{i})\geq i+2$, and
iii) For $i=1,2,\ldots,n-2$, $N^{+}_{R}(v_{i})=\{v_{min(v_{i})},v_{min(v_{i})+1},\ldots,v_{max(v_{i})}\}$.
Proof.
i) Edges $v_{1}v_{2},v_{1}v_{3},v_{n-2}v_{n}$ and $v_{n-1}v_{n}$ do not cross any edge of $T_{n}$. Therefore, they all must lie in $R$.
ii) Clearly $min_{R}(v_{i+1})\geq max_{R}(v_{i})$ for $i=1,2,\ldots,n-2$, otherwise $v_{i+1}v_{min_{R}(v_{i+1})}$ and $v_{i}v_{max_{R}(v_{i})}$ would be edges of $R$ that cross each other. Suppose $min_{R}(v_{t+1})>max_{R}(v_{t})$ for some integer $1\leq t\leq n-2$; in particular $v_{t}v_{min_{R}(v_{t+1})}$ is not an edge of $R$. Since $R$ is a maximal plane subgraph, edge $v_{t}v_{min_{R}(v_{t+1})}$ must cross some edge $v_{q}v_{r}$ of $R$. If $q<t<min_{R}(v_{t+1})<r$, then $q<t<max_{R}(v_{t})<r$ which imples that edges $v_{q}v_{r}$ and $v_{t}v_{max_{R}(v_{t})}$ of $R$ cross each other, a contradiction. Suppose now $t<q<r<min_{R}(v_{t+1})$. Notice that $q\neq t+1$, otherwise $v_{t+1}v_{r}$ is an edge of $R$ with $t+1<r<min_{R}(v_{t+1})$, which is a contradiction. Then $t+1<q<r<min_{R}(v_{t+1})$, in which case edges $v_{q}v_{r}$ and $v_{t+1}v_{min_{R}(v_{t+1})}$ cross each other in $R$, once again a contradiction.
By i), $max_{R}(v_{1})\geq 3$. Suppose $max_{R}(v_{j})\geq j+2$ for $j=1,2,\ldots,t$ and $max_{R}(v_{t+1})\leq t+2$ for some integer $1\leq t<n-2$. Let $l=max\{max_{R}(v_{t}),t+3\}$. Notice that $v_{t+1}v_{l}$ is not an edge of $R$ since $max_{R}(v_{t+1})\leq t+2<l$. As $R$ is a maximal plane subgraph of $T_{n}$, there is an edge $v_{q}v_{r}$ of $R$ that crosses $v_{t+1}v_{l}$. If $q<t+1<l<r$, then either $v_{q}v_{r}$ is an edge of $R$ with $q=t$ and $max_{R(v_{t})}<r$ which is a contradiction or $q<t<l<r$ which implies that edges $v_{q}v_{r}$ and $v_{t}v_{max_{R}(v_{t})}$ cross each other in $R$, again a contradiction. Therefore $max_{R}(v_{i})\geq i+2$ for $i=1,2,\ldots,n-2$.
iii) Let $t=1,2,\ldots,n-2$ and suppose $v_{t}v_{j}\notin N^{+}_{R}(v_{t})$ for some $min_{R}(v_{t})<j<max_{R}(v_{t})$. Since $R$ is a maximal plane subgraph of $T_{n}$, there is an edge $v_{q}v_{r}$ of $R$ that crosses the edge $v_{t}v_{j}$. If $q<t<j<r$, then $q<t<min_{R}(v_{t})<r$ and therefore $v_{q}v_{r}$ would cross the edge $v_{t}v_{min(v_{t})}$ of $R$, which is not possible. Then $t<q<r<max_{R}(v_{t})$ which implies $v_{q}v_{r}$ crosses the edge $v_{t}v_{max(v_{t})}$ of $R$. Also a contradiction.
Notice that for any maximal plane subgraph $R$ of $T_{n}$ with $n\geq 3$, Lemma 2 implies $N^{+}_{R}(v_{i})\neq\emptyset$ for $i=1,2,\ldots,n-1$. For any pair $R$, $Q$ of maximal plane subgraphs of $T_{n}$ let
[…]
and let m(R,Q)=n1k(R,Q)m(R,Q)=n-1-k(R,Q).
Let $F$ be a set of pairwise non-crossing edges of $T_{n}$. We denote by $MP(T_{n},F)$ the subgraph of $MP(T_{n})$ induced by the maximal plane subgraphs of $T_{n}$ that contain all edges in $F$. The following result will be used, together with Theorem 1, in the proof of our main result.
Theorem 3.
For each integer $n\geq 3$ and each set $F$ of pairwise non-crossing edges of $T_{n}$, the graph $MP(T_{n},F)$ is connected.
Proof.
Let $F$ be a set of pairwise non-crossing edges of $T_{n}$ and $R$ and $Q$ be maximal plane subgraphs of $T_{n}$ with $F\subset E(R)\cap E(Q)$. We shall proof by induction on $m(R,Q)$ that $MP(T_{n},F)$ contains a path joining $R$ and $Q$.
If $m(R,Q)=0$, then k(R,Q)=n1k(R,Q)=n-1. This implies $N_{R}^{+}(v_{i})=N_{Q}^{+}(v_{i})$ for $i=1,2,\ldots,n-1$ and therefore, $R=Q$. We proceed assuming that for some integer $t\geq 0$ the result holds whenever $R^{\prime}$ and $Q^{\prime}$ are maximal planar subgraphs of $T_{n}$ with $F\subset E(R^{\prime})\cap E(Q^{\prime})$ and $m(R^{\prime},Q^{\prime})\leq t$.
Let $R$ and $Q$ be maximal planar subgraphs of $T_{n}$ with $F\subset E(R)\cap E(Q)$ and $m(R,Q)=t+1$. Let $k=k(R,Q)$.
Then, $N_{R}^{+}(v_{i})=N_{Q}^{+}(v_{i})\text{ for }i\leq k$ and $N_{R}^{+}(v_{k+1})\neq N_{Q}^{+}(v_{k+1})$; without loss of generality we assume $d_{R}^{+}(v_{k+1})>d_{Q}^{+}(v_{k+1})$. Notice that by Lemma 2, $min_{R}(v_{k+1})=max_{R}(v_{k})=max_{Q}(v_{k})=min_{Q}(v_{k+1})$. Then $max_{R}(v_{k+1})>max_{Q}(v_{k+1})$, and therefore $v_{k+1}v_{max_{R}(v_{k+1})}$ is an edge of $R$ which is not an edge of $Q$. Also by Lemma 2, $max_{Q}(v_{k+1})\geq k+3$; this implies $max_{R}(v_{k+1})\geq k+4$ and therefore $v_{k+2}v_{max_{R}(v_{k+1})-1}$ is an edge of $T_{n}$. Notice that $v_{k+2}v_{max_{R}(v_{k+1})-1}$ is not an edge of $R$ since it crosses the edge $v_{k+1}v_{max_{R}(v_{k+1})}$ which does lie in $R$. Let
[…]
Figure 2: Graphs $R$ (above) and $S_{1}$ (bellow).
We claim $S_{1}$ is a maximal plane subgraph of $T_{n}$ adjacent to $R$ in $MP(T_{n},F)$. To prove this, assume first that $S_{1}$ is not a plane graph, then the added edge $v_{k+2}v_{max_{R}(v_{k+1})-1}$ must cross an edge $v_{q}v_{r}$ of $R$, other than the deleted edge $v_{k+1}v_{max_{R}(v_{k+1})}$. If $k+2<q<r<max_{R}(v_{k+1})-1$, then $v_{q}v_{r}$ crosses the edge $v_{k+1}v_{max_{R}(v_{k+1})}$ in $R$ which is not possible. Suppose now $q<k+2<max_{R}(v_{k+1})-1<r$. If $q=k+1$, then $v_{q}v_{r}$ is an edge of $R$ with $q=k+1$ and $r\geq max_{R}(v_{k+1})$; this implies $v_{q}v_{r}$ is the edge $v_{k+1}v_{max_{R}(v_{k+1})}$ which is not an edge of $S_{1}$, therefore $q<k+1$. In this case notice that $v_{q}v_{r}$ crosses the edge $v_{k+1}v_{max_{R}(v_{k+1})-1}$ which is an edge of $R$ by Lemma 2, since $min_{R}(v_{k+1})=min_{Q}(v_{k+1})\leq max_{Q}(v_{k+1})<max_{R}(v_{k+1})$. Again this is not possible.
Next, assume there is an edge $v_{q}v_{r}$ of $T_{n}$, not in $S_{1}$, such that $S_{1}+v_{q}v_{r}$ is a plane graph. Since $R$ is a maximal plane subgraph of $T_{n}$, edge $v_{q}v_{r}$ must cross the edge $v_{k+1}v_{max_{R}(v_{k+1})}$ which is the only edge of $R$ not in $S_{1}$. If $q<k+1<max_{R}(v_{k+1})<r$, then $v_{q}v_{r}$ crosses the edge $v_{k+2}v_{max_{R}(v_{k+1})-1}$ of $S_{1}$; and if $k+1<q<r<max_{R}(v_{k+1})$, then $v_{q}v_{r}$ would cross the edge $v_{k+2}v_{max_{R}(v_{k+1})}$ which is an edge of $S_{1}$ since, by Lemma 2, $min_{R}(v_{k+2})=max_{R}(v_{k+1})$. This ends the proof of our claim.
Notice that $F\subset E(S_{1})$ since $F\subset E(R)\cap E(Q)$ and $v_{k+1}v_{max_{R}(v_{k+1})}\in E(R)\setminus E(Q)$; therefore $R$ and $S_{1}$ are adjacent in $MP(T_{n},F)$.
If $N_{S_{1}}^{+}(v_{k+1})=N_{Q}^{+}(v_{k+1})$, then $k(S_{1},Q)\geq k+1$ and $m(S_{1},Q)\leq t$. By the induction hypothesis $MP(T_{n},F)$ contains a path joining $S_{1}$ and $Q$ which, together with the fact that $R$ and $S_{1}$ are adjacent in $MP(T_{n},F)$, gives a path in $MP(T_{n},F)$ that joins $R$ and $Q$.
If $N_{S_{1}}^{+}(v_{k+1})\neq N_{Q}^{+}(v_{k+1})$, then $max_{S_{1}}(v_{k+1})>max_{Q}(v_{k+1})$. We iterate the process defining, one at a time, maximal plane subgraphs $S_{2},S_{3},\ldots,S_{l}$ of $T_{n}$, where $l=max_{R}(v_{k+1})-max_{Q}(v_{k+1})$.
$S_{i+1}=(S_{i}-v_{k+1}v_{max_{S_{i}}(v_{k+1})})+v_{k+2}v_{max_{S_{i}}(v_{k+1})-1}.$
Since $F\subset E(R)\cap E(Q)$ and $v_{k+1}v_{max_{s_{i}}(v_{k+1})}=v_{k+1}v_{max_{R}(v_{k+1})-i}\in E(R)\setminus E(Q)$ for $i=1,2,\ldots l-1$, $v_{k+1}v_{max_{s_{i}}(v_{k+1})}\notin F$ for $i=1,2,\ldots l-1$. This implies $F\subset E(S_{i})$ for $i=1,2,\ldots,l$. Therefore $R,S_{1},S_{2},\ldots,S_{l}$ is a path in $MP(T_{n},F)$ joining $R$ and $S_{l}$.
Notice that $N_{S_{i+1}}^{+}(v_{k+1})=N_{S_{i}}^{+}(v_{k+1})\setminus\{v_{k+1}v_{max_{S_{i}}(v_{k+1})}\}$ for $i=1,2,\ldots,l-1$, therefore $N_{S_{l}}^{+}(v_{k+1})=N_{Q}^{+}(v_{k+1})$. Then $k(S_{l},Q)\geq k+1$ and $m(S_{l},Q)\leq t$. Again by the induction hypothesis, $MP(T_{n},F)$ contains a path joining $S_{l}$ and $Q$ which, together with the path $R,S_{1},S_{2},\ldots,S_{l}$, forms a path in $MP(T_{n},F)$ that joins $R$ and $Q$.
3 Main result
As shown in Fig. 3, not all maximal plane subgraphs of $T_{2m}$ contain perfect matchings. Let $MP_{\cal{M}}(T_{2m})$ denote the subgraph of $MP(T_{2m})$ induced by the set of maximal plane subgraphs of $T_{2m}$ that contain at least one perfect matching.
Figure 3: A maximal plane subgraph ot $T_{6}$ with no perfect matchings.
Theorem 4.
For each positive integer $m$, the graph $MP_{\cal{M}}(T_{2m})$ is connected.
Proof.
Let $S$ and $R$ be maximal plane subgraphs of $T_{2m}$, each containing at least one perfect matching, and let $M(S)$ and $M(R)$ be perfect matchings of $S$ and $R$, respectively. By Theorem 1, the graph $\mathcal{M}\left(T_{2m}\right)$ is connected and therefore contains a path $M(S)=M_{0},M_{1},\ldots,M_{k}=M(R)$ joining $M(S)$ and $M(R)$.
By definition, for $i=1,2,\ldots,k$, $M_{i}$ is obtained from $M_{i-1}$ by exchanging the edges of $M_{i}$ and $M_{i-1}$ of an alternating non-crossing cycle with four edges. Therefore, for $i=1,2,\ldots,k$, $M_{i-1}\cup M_{i}$ is a set of pairwise non-crossing edges of $T_{2m}$ that can be extended to a maximal plane subgraph $S_{i}$ of $T_{2m}$.
Let $S_{0}=S$ and $S_{k+1}=R$. Notice that for $i=0,1,\ldots,k$, $S_{i}$ and $S_{i+1}$ are maximal plane graphs of $T_{2m}$ that contain $M_{i}$. By Theorem 3, the graph $MP(T_{2m},M_{i})$ contains a path joining $S_{i}$ and $S_{i+1}$. Clearly $MP(T_{2m},M_{i})$ is a subgraph of $MP_{\cal{M}}(T_{2m})$ for each $i=0,1,\ldots,k$. Therefore the graph $MP_{\cal{M}}(T_{2m})$ contains a path joining $S$ and $R$. ∎